# Efficient Trustworthy Distillation

### v1

@@ Abstract
Large language models (LLMs) have become increasingly prevalent in our daily lives, leading to an expectation for LLMs to be trustworthy —- both accurate and well-calibrated (the prediction confidence should align with its ground truth correctness likelihood). Nowadays, fine-tuning has become the most popular method for adapting a model to practical usage by significantly increasing accuracy on downstream tasks. Despite the great accuracy it achieves, we found fine-tuning is still far away from satisfactory trustworthiness due to " tuning-induced mis-calibration". In this paper, we delve deeply into why and how mis-calibration exists in fine-tuned models, and how distillation can alleviate the issue. Then we further propose a brand new method named EFfIcient TRustworthy DiSTillation (FIRST), which utilizes a small portion of teacher’s knowledge to obtain a reliable language model in a cost-efficient way. Specifically, we identify the "concentrated knowledge" phenomenon during distillation, which can significantly reduce the computational burden. Then we apply a "trustworthy maximization" process to optimize the utilization of this small portion of concentrated knowledge before transferring it to the student. Experimental results demonstrate the effectiveness of our method, where better accuracy (+2.3%) and less mis-calibration (-10%) are achieved on average across both in-domain and out-of-domain scenarios, indicating better trustworthiness.
@@
Kashun Shum♡∗,   Minrui Xu♡∗,   Jianshu Zhang♠∗,   Zixin Chen,   Shizhe Diao   Hanze Dong,   Jipeng Zhang ,   Muhammad Omer Raza The Hong Kong University of Science and Technology, Wuhan University Purdue University {ksshumab, mxubh}@connect.ust.hk
@@ 1 Introduction
With the rapid development of large language models (LLMs), many powerful models have been deployed into our daily lives for practical usage to help us make decisions (yao2023react; sha2023languagempc; zhao2024expel). This makes it urgent for us to know to what extent we can trust the outputs of the models. Calibration is one of the most important indicators beyond accuracy which provides a confidence measure to the model’s predictions (guo2017calibration; hsieh-etal-2023-distilling). In LLMs, confidence is exactly the probability for each generated token. Therefore, a well-calibrated model should align its prediction confidence with its ground-truth correctness likelihood. As an example, recent hallucination detection methods rely on model prediction confidence as a significant indicator of potential hallucination (zhang-etal-2023-enhancing-uncertainty; Varshney2023ASI). If the model is incapable of giving accurate confidence levels, people may fail to detect hallucinations due to the model’s over-confidence, or people may falsely identify hallucinations due to the model’s under-confidence. Mis-calibration brings significant challenges for the deployment of LLMs in real-world applications.
Currently, there are two methods to obtain a language model for practical usage. First, fine-tuning, which fine-tunes pre-trained LLMs on specific datasets by matching each token entry with a target ground truth token. Although fine-tuning can consistently improve performance on downstream tasks (dodge2020finetuning; sun2020finetune; ziegler2020finetuning), we identify that the model obtained in this way exhibits a nature of "tuning-induced miscalibration". Second, distillation-based methods transfer knowledge (e.g., soft labels) from larger LLMs to smaller models (gu2023knowledge). Although distillation shows better calibration than fine-tuning as it matches each token entry with a probability distribution instead of a hard label, we find it is still biased because of the mis-calibration nature of teacher models. In addition, distillation faces the challenge of determining the optimal amount of knowledge to transfer. Transferring all the teacher’s knowledge leads to high computational costs while transferring too little knowledge results in poor accuracy. Therefore, it is crucial to balance between trustworthiness (accuracy and well-calibration) and efficiency for distillation-based methods.
@@
We first validate our method in in-domain scenarios, discovering that the models obtained by FIRST achieve excellent accuracy, even with the use of a relatively small amount of top-5 knowledge and the "trustworthy maximization" process can significantly enhance these models’ calibration ability. Furthermore, we test our approach in out-of-domain settings, demonstrating that models obtained by FIRST still exhibit the best trustworthiness and hold generalization ability. This indicates that FIRST enables smaller models to genuinely learn the capability of being trustworthy, rather than being confined to in-domain scenarios.
@@ 2.1 Trustworthy Models
The current evaluation of LLMs predominantly focuses on accuracy, overlooking whether the models truly know the answer or are merely guessing (i.e. trustworthy). Recent works (sun2024trustllm; steyvers2024calibration) have demonstrated that accurate LLMs may not necessarily be "trustworthy" due to a significant calibration gap, so-called mis-calibration. This gap prevents us from trusting the output of the models, and it can further cause LLMs to generate harmful content, especially when subjected to adversarial attacks or jailbreak prompts (mo2024trustworthy; yao2024fuzzllm). Our work further reveals how mis-calibration exists in different tuning methods and proposes a new trustworthy evaluation metric that covers both accuracy and calibration.
To achieve a well-calibrated LLM, recent work shows soft-label distillation shows better calibration ability (gu2023knowledge). However, it still suffers from biased labels due to the mis-calibration nature of the fine-tuned teacher model. Our work is an improvement on this line of work by applying "concentrated knowledge" and "trustworthy maximization", leading to better accuracy, efficiency, and trustworthy.
@@ 2.2 Knowledge Distillation
Black-box Distillation refers to distillation from models that we are unable to access the weight and prediction logits such as PaLM (chowdhery2022palm). Recent studies have attempted to distill reasoning ability from GPT (ho-etal-2023-large; shridhar-etal-2023-distilling) or some emergent ability such as chain-of-thought (hsieh-etal-2023-distilling; Li2023SymbolicCD). However, these methods may still be categorized as the genre of data-augmentation-and-then-fine-tuning approaches.
@@
White-box Distillation means the teacher models are either fully open-sourced such as Llama (touvron2023llama) or they can return partial probability distribution of the generated tokens, such as code-davinci-002. Instead of the hard token fine-tuning, white-box distillation typically uses more fine-grained signals by matching a distribution between teachers and students (gu2023knowledge; latif2023knowledge; agarwal2024onpolicy). Further, in the field of white-box distillation, there are two different ways: online distillation and offline distillation. Onlin distillation (gu2023knowledge; zhou2023distillspec) needs to keep both the teacher model and the student model on the GPU simultaneously during training. On the other hand, offline distillation typically involves obtaining knowledge from the teacher model beforehand. Our work is an extension of white-box offline distillation and focuses on how white-box offline distillation can be improved in terms of trustworthiness by re-calibrating the teacher distribution.
@@ 3.3 Expected Calibration Error
To measure calibration in the context of LLMs, we adapt the expected calibration error (ECE) to the free-text generation task by treating the generation of a single token as a classification task. In this adaptation, we restrict the model to generate only one token from a set of candidate choices (e.g., A/B/C/D). For each token, we obtain the highest probability choice using $\mathop{\arg\max}_{i\in C}P(i)$, where $C$ represents the set of candidates. The probability of the chosen token is taken as the predicted confidence, and we calculate the accuracy by comparing the predicted choice to the ground truth. Then we utilize a total $M$ probability interval as bins and categorize each chosen token into $m$-th bin according to the predicted confidence. The ECE can be computed as follows:
@@ 3.4 Trustworthy Score
In evaluating the trustworthiness of a model, it is essential to consider both high accuracy and effective calibration. Existing benchmarks primarily focus on accuracy, assuming that higher accuracy implies greater trustworthiness. However, our discovery of the widespread issue of "tuning-induced mis-calibration" has highlighted the inadequacy of relying solely on accuracy for a comprehensive evaluation of model reliability. To address this limitation, we propose Trust Score metric to quantify a model’s trustworthiness, which quantifies the trustworthiness of a model by considering two key aspects: its ability to provide accurate answers (measured by $Acc$) and its capacity to align predicted confidences with actual accuracies (measured by $ECE$). The Trust Score is defined as follows:
@@ 4 Efficient Trustworthy Distillation
Figure 4: (a) The Trustworthy Maximization Pipeline. We firstly randomly select a validation set and fine-tuning our teacher models. Then based on the validation set, we generate top-5 probabilities of all tokens and run a grid search to select the optimal temperature. (b) The Efficient Trustworthy Distillation Pipeline. We obtained a well-calibrated student model based on the optimal temperature from the trustworthy maximization step.
In this section, we introduce eFfIcient tRustworthy disTillation (FIRST), which can be divided into three parts. Firstly, we select Top-5 tokens as knowledge for transfer (Efficient Knowledge Selection) in Sec.4.1. Then, we adjust the knowledge for trustworthiness to ensure that the subsequent smaller models can maximize its utility (Knowledge Trustworthy Maximization) in Sec.4.2. Finally, we describe the learning process of the student model (Knowledge Matching) in Sec.4.3.
@@ 4.2 Trustworthy Maximization
Label Smoothing.
We first attempted to address tuning-induced mis-calibration" by applying a smoothing coefficient, denoted as $\delta$, to mitigate the teacher model’s over-confidence in its top-1 token predictions while alleviating under-confidence in other predicted tokens as follows:
@@
Temperature Scaling.
Subsequently, we explore another approach using a temperature scaling technique to re-calibrate the probabilities:
@@
This method offers several advantages. First, it allows for a more fine-grained adjustment of the probability distribution by controlling the temperature scaling parameter $c$, which can be optimized to achieve the lowest ECE values. Second, unlike label smoothing, temperature scaling can effectively balance the confidence levels of both top-1 and subsequent tokens, reducing both over-confidence and under-confidence issues. This results in a more consistent and reliable calibration across all tokens, thereby enhancing the overall trustworthiness of the knowledge. Additionally, we find that selecting the optimal $c$ parameter on the validation set to maximize the knowledge significantly enhances the effectiveness of transferring trustworthy knowledge. The knowledge processed by using this $c$ yields the best results for the student model (detailed in Sec. §5.5). Due to the low cost of selecting $c$ on the validation set, we can tailor different $c$ values for different tasks. This demonstrates "temperature scaling" excellent scalability and flexibility.
@@ 5.1 Experimental Settings
Our experiments focus on both In-Domain and Out-of-Domain settings to ensure generalization abilities. In the In-Domain setting, we utilize CommonsenseQA (CSQA) (talmor2019commonsenseqa) and BoolQ (clark2019boolq) for both training and testing. In the Out-of-Domain setting, we fine-tune and distill smaller models on a commonly used instruction-following dataset, Alpaca (alpaca), while, testing the models’ performance over unseen task CommonsenseQA (CSQA) and OpenBook QA (OBQA) (mihaylov-2018-suit). This approach allows us to assess the generalization abilities of the smaller models on unseen tasks, simulating real-world scenarios where these models need to perform on unfamiliar tasks.
To ensure the practicality of our approach, we select three widely used model families for our experiments: Llama-1 (touvron2023llama), Llama-2 (touvron2023llama2), and OpenLlama (openlm2023openllama). In our experiments, we test four types of smaller models obtained through different methods:
@@
4) Distill 7B w/ LS: We also explore the use of label smoothing (LS, see Eq. 3) to show why we ultimately adopt TS over LS in "knowledge trustworthy maximization". In the latter experiments, we pick up the popular smoothing coefficient 0.1 follow previous works (müller2020does). Additionally, we also provide the performance of Teacher models. For further implementation details, please refer to the Appendix.
@@ 6 Conclusion
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. We address the critical issue of catastrophic mis-calibration in current training pipelines (supervised fine-tuning and knowledge distillation) and propose a pipeline to efficiently obtain a more trustworthy model. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ 8 Limitations
It is shown that our efficient trustworthy distillation (FIRST) demonstrates superior calibration ability and performance over direct distillation and standard fine-tuning methods. However, despite these exciting results, there are still some limitations to our current work, as well as potential opportunities for future research.
@@
: Due to the resource limitation, our largest teacher model is Llama 33B which is not very large but already achieving exciting results by distillation to a 7B student model. We expect that employing a large teacher model such as 70B can lead to better calibration ability and performance since a large model learns a better distribution. However, we are unable to explore how very large teachers perform due to resource limitations.
@@
Standard Fine-tuning Direct Distillation FIRST
Question Which city is farther north, Oslo or Helsinki?
Correct Answer Helsinki
Generated Oslo is farther north than Helsinki. Oslo is farther north than Helsinki. Oslo is farther north than Helsinki.
[…]
Question Is Donald Trump a Neo-con American politician and businessman for the Republicans, with a long and varied career?
Correct Answer No
Generated Yes. Yes. Yes.
[…]
Question If I want to visit Beijing in spring, when should I go? Answer Choices: (a) June (b) July (c) August (d) September (e) October
Correct Answer None
Generated (c). (d). (d).
[…]
Table 2: A case study on how fine-tuned model and direct distilled model tend to over-confident on the wrong answer with high confidence. While FIRST though outputs a wrong answer, it produces low confidence to show its uncertainty.
@@ A.1 Implementation Details
We train our models on 8 GPU (RTX A6000 48G) using the Adam optimizer with beta set to be [0.9, 0.999] and epsilon fixed to be 1e-6 and cosine annealing scheduler with a warm-up ratio of 0.03. For fine-tuning, we utilize LMFlow (diao2023lmflow) package to obtain a well fine-tuned model by a standard 3-epoch training and control the batch size to be 32 on each GPU and the learning rate for teacher models to be 2e-5. Finally, for distillation, the batch size is set to 32 on each GPU and we train our model for 3 epochs, the last checkpoint is used for evaluation since it has the best performance.

### v2

@@ Abstract
Large language models (LLMs) have become increasingly prevalent in our daily lives, leading to an expectation for LLMs to be trustworthy — both accurate and well-calibrated (the prediction confidence should align with its ground truth correctness likelihood). Nowadays, fine-tuning has become the most popular method for adapting a model to practical usage by significantly increasing accuracy on downstream tasks. Despite the great accuracy it achieves, we found fine-tuning is still far away from satisfactory trustworthiness due to "tuning-induced mis-calibration". In this paper, we delve deeply into why and how mis-calibration exists in fine-tuned models, and how distillation can alleviate the issue. Then we further propose a brand new method named EFfIcient TRustworthy DiSTillation (FIRST), which utilizes a small portion of teacher’s knowledge to obtain a reliable language model in a cost-efficient way. Specifically, we identify the "concentrated knowledge" phenomenon during distillation, which can significantly reduce the computational burden. Then we apply a "trustworthy maximization" process to optimize the utilization of this small portion of concentrated knowledge before transferring it to the student. Experimental results demonstrate the effectiveness of our method, where better accuracy (+2.3%) and less mis-calibration (-10%) are achieved on average across both in-domain and out-of-domain scenarios, indicating better trustworthiness. [^1] The code is available at https://github.com/SHUMKASHUN/FIRST.
@@
Kashun Shum♡∗,   Minrui Xu♡∗,   Jianshu Zhang♠∗,   Zixin Chen,   Shizhe Diao   Hanze Dong,   Jipeng Zhang ,   Muhammad Omer Raza The Hong Kong University of Science and Technology, Wuhan University NVIDIA, Purdue University {ksshumab, mxubh}@connect.ust.hk
@@ 1 Introduction
With the rapid development of large language models (LLMs), many powerful models have been deployed into our daily lives for practical usage to help us make decisions (Yao et al., 2023; Sha et al., 2023; Zhao et al., 2024). This makes it urgent for us to know to what extent we can trust the outputs of the models. Calibration is one of the most important indicators beyond accuracy, which provides a confidence measure to the model’s predictions (Guo et al., 2017; Hsieh et al., 2023). In LLMs, confidence is exactly the probability for each generated token. Therefore, a well-calibrated model should align its prediction confidence with its ground-truth correctness likelihood as shown in Figure 1. As an example, recent hallucination detection methods rely on model prediction confidence as a significant indicator of potential hallucination (Zhang et al., 2023; Varshney et al., 2023). If the model is incapable of giving accurate confidence levels, people may fail to detect hallucinations due to the model’s over-confidence, or people may falsely identify hallucinations due to the model’s under-confidence. Mis-calibration brings significant challenges for the deployment of LLMs in real-world applications.
Currently, there are two methods to obtain a language model for practical usage. First, fine-tuning, which fine-tunes pre-trained LLMs on specific datasets by matching each token entry with a target ground truth token. Although fine-tuning can consistently improve performance on downstream tasks (Dodge et al., 2020; Sun et al., 2020; Ziegler et al., 2020), we identify that the model obtained in this way exhibits a nature of "tuning-induced mis-calibration". Second, distillation-based methods transfer knowledge (e.g., soft labels) from larger LLMs to smaller models (Gu et al., 2023). Although distillation shows better calibration than fine-tuning as it matches each token entry with a probability distribution instead of a hard label, we find it is still biased because of the mis-calibration nature of teacher models. In addition, distillation faces the challenge of determining the optimal amount of knowledge to transfer. Transferring all the teacher’s knowledge leads to high computational costs while transferring too little knowledge results in poor accuracy. Therefore, it is crucial to balance between trustworthiness (accuracy and well-calibration) and efficiency for distillation-based methods.
@@
We first validate our method in in-domain scenarios, discovering that the models obtained by FIRST achieve excellent accuracy, even with the use of a relatively small amount of top-5 knowledge and the "trustworthy maximization" process can significantly enhance these models’ robustness to mis-calibration. Furthermore, we test our approach in out-of-domain settings, demonstrating that models obtained by FIRST still exhibit the best trustworthiness and hold generalization ability. This indicates that FIRST enables smaller models to genuinely learn the capability of being trustworthy, rather than being confined to in-domain scenarios.
@@ 2.1 Trustworthy Models
The current evaluation of LLMs predominantly focuses on accuracy, overlooking whether the models truly know the answer or are merely guessing (i.e. trustworthy). Recent works (Sun et al., 2024; Steyvers et al., 2024) have demonstrated that accurate LLMs may not necessarily be "trustworthy" due to a significant calibration gap, so-called mis-calibration. This gap prevents us from trusting the output of the models, and it can further cause LLMs to generate harmful content, especially when subjected to adversarial attacks or jailbreak prompts (Mo et al., 2024; Yao et al., 2024). Our work further reveals how mis-calibration exists in different tuning methods and proposes a new trustworthy evaluation metric that covers both accuracy and calibration.
To achieve a well-calibrated LLM, recent work shows soft-label distillation shows better calibration ability (Gu et al., 2023). However, it still suffers from biased labels due to the mis-calibration nature of the fine-tuned teacher model. Our work is an improvement on this line of work by applying "concentrated knowledge" and "trustworthy maximization", leading to better accuracy, efficiency, and trustworthy.
@@ 2.2 Knowledge Distillation
Black-box Distillation refers to distillation from models that we are unable to access the weight and prediction logits such as PaLM (Chowdhery et al., 2022). Recent studies have attempted to distill reasoning ability from GPT (Ho et al., 2023; Shridhar et al., 2023) or some emergent ability such as chain-of-thought (Hsieh et al., 2023; Li et al., 2023). However, these methods may still be categorized as the genre of data-augmentation-and-then-fine-tuning approaches.
@@
White-box Distillation means the teacher models are either fully open-sourced such as Llama (Touvron et al., 2023a) or they can return partial probability distribution of the generated tokens, such as code-davinci-002. Instead of the hard token fine-tuning, white-box distillation typically uses more fine-grained signals by matching a distribution between teachers and students (Gu et al., 2023; Latif et al., 2023; Agarwal et al., 2024). Further, in the field of white-box distillation, there are two different ways: online distillation and offline distillation. Online distillation (Gu et al., 2023; Zhou et al., 2023) needs to keep both the teacher model and the student model on the GPU simultaneously during training. On the other hand, offline distillation typically involves obtaining knowledge from the teacher model beforehand. Our work is an extension of white-box offline distillation and focuses on how white-box offline distillation can be improved in terms of trustworthiness by re-calibrating the teacher distribution.
@@ 3.3 Expected Calibration Error
To measure calibration in the context of LLMs, we adapt the expected calibration error (ECE) to the free-text generation task by treating the generation of a single token as a classification task. In this adaptation, we restrict the model to generate only one token from a set of candidate choices (e.g., A/B/C/D). For each token, we obtain the highest probability choice using $\mathop{\arg\max}_{i\in C}P(i)$, where $C$ represents the set of candidates. The probability of the chosen token is taken as the predicted confidence, and we calculate the accuracy by comparing the predicted choice to the ground truth. Then we utilize a total $M$ probability intervals as bins and categorize each chosen token into $m$-th bin according to the predicted confidence. The ECE (Guo et al., 2017) can be computed as follows:
@@ 3.4 Trustworthy Score
When evaluating the trustworthiness of a model, it is essential to consider both high accuracy and effective calibration. Existing benchmarks primarily focus on accuracy, assuming that higher accuracy implies greater trustworthiness. However, our discovery of the widespread issue of "tuning-induced mis-calibration" has highlighted the inadequacy of relying solely on accuracy for a comprehensive evaluation of model trustworthiness. To address this limitation, we propose Trust Score metric to quantify a model’s trustworthiness, which considers two key aspects: its ability to provide accurate answers (measured by $Acc$) and its capacity to align predicted confidences with actual accuracies (measured by $ECE$). The Trust Score is defined as follows:
@@ 4 Efficient Trustworthy Distillation
Figure 4: (a) The Trustworthy Maximization Step: we first fine-tune our the teacher model and then generate top-5 probabilities of all tokens and run a grid search to select the optimal temperature based on the validation set. (b) The overall Efficient Trustworthy Distillation Pipeline: based on the selected optimal temperature from (a), we obtain a well-calibrated student model by knowledge matching between student’s knowledge and the portion of teacher knowledge.
In this section, we introduce eFfIcient tRustworthy disTillation (FIRST), which can be divided into three parts. Firstly, we select top-5 tokens as knowledge for transfer (Efficient Knowledge Selection) in Sec. §4.1. Then, we adjust the knowledge for trustworthiness to ensure that the subsequent smaller models can maximize its utility (Knowledge Trustworthy Maximization) in Sec. §4.2. Finally, we describe the learning process of the student model (Knowledge Matching) in Sec. §4.3. The overall pipeline is shown in Figrue 4.
@@ 4.2 Trustworthy Maximization
Label Smoothing:
Similar to Müller et al. (2019), we first attempted to address tuning-induced mis-calibration" by applying a smoothing coefficient, denoted as $\delta$, to mitigate the teacher model’s over-confidence in its top-1 token predictions while alleviating under-confidence in other predicted tokens as follows:
@@
Temperature Scaling:
Subsequently, we explore another approach using a temperature scaling technique (Guo et al., 2017) to re-calibrate the probabilities:
@@
This method offers several advantages. First, it allows for a more fine-grained adjustment of the probability distribution by controlling the temperature scaling parameter $c$, which can be optimized to achieve the lowest ECE values. Second, unlike label smoothing, temperature scaling can effectively balance the confidence levels of both top-1 and subsequent tokens, reducing both over-confidence and under-confidence issues by preserving token rankings and avoiding transition between under-confidence and over-confidence.
This results in a more consistent and reliable calibration across all tokens, thereby enhancing the overall trustworthiness of the knowledge. Additionally, we find that selecting the optimal $c$ parameter on the validation set to maximize the knowledge can significantly enhance the effectiveness of transferring trustworthy knowledge. The knowledge processed by using this $c$ yields the best results for the student model (detailed in Sec. §5.5). Due to the low cost of selecting $c$ on the validation set, we can tailor different $c$ values for different tasks. This demonstrates "temperature scaling" excellent scalability and flexibility.
@@ 5.1 Experimental Settings
Our experiments focus on both In-Domain and Out-of-Domain settings to ensure generalization abilities. In the In-Domain setting, we utilize CommonsenseQA (CSQA) (Talmor et al., 2019) and BoolQ (Clark et al., 2019) for both training and testing. In the Out-of-Domain setting, we fine-tune and distill smaller models on a commonly used instruction-following dataset, Alpaca (Taori et al., 2023), while, testing the models’ performance over unseen task CommonsenseQA (CSQA) and OpenBook QA (OBQA) (Mihaylov et al., 2018). This approach allows us to assess the generalization abilities of the smaller models on unseen tasks, simulating real-world scenarios where these models need to perform on unfamiliar tasks.
To ensure the practicality of our approach, we select three widely used model families for our experiments: Llama-1 (Touvron et al., 2023a), Llama-2 (Touvron et al., 2023b), and OpenLlama (Geng and Liu, 2023). In our experiments, we test four types of smaller models obtained through different methods:
@@
4) Distill 7B w/ LS: We also explore the use of label smoothing (LS, see Eq. 3) to show why we ultimately adopt TS over LS in "knowledge trustworthy maximization". In the latter experiments, we pick up the popular smoothing coefficient 0.1 follow previous works (Müller et al., 2020).
Additionally, we also provide the performance of Teacher models. For further implementation details, please refer to the Appendix A.
@@ 7 Limitations
It is shown that our efficient trustworthy distillation (FIRST) demonstrates superior calibration ability and performance over direct distillation and standard fine-tuning methods. However, despite these exciting results, there are still some limitations in our current work, as well as potential opportunities for future research.
@@
: Due to the resource limitation, our largest teacher model is Llama 33B, which is not very large but already achieving exciting results by distillation to a 7B student model. We expect that employing a large teacher model such as 70B can lead to better calibration ability and performance since a larger model learns a better distribution. However, we are unable to explore how very large teachers perform due to resource limitations.
@@ A.1 Implementation Details
We train our models on 8 GPU (RTX A6000 48G) using the Adam optimizer with beta set to be [0.9, 0.999] and epsilon fixed to be 1e-6 and cosine annealing scheduler with a warm-up ratio of 0.03. For fine-tuning, we utilize LMFlow (Diao et al., 2023) package to obtain a well fine-tuned model by a standard 3-epoch training and control the batch size to be 32 on each GPU and the learning rate for teacher models to be 2e-5. Finally, for distillation, the batch size is set to 32 on each GPU and we train our model for 3 epochs, the last checkpoint is used for evaluation since it has the best performance.